# Generative Autoregressive Transformers for Model-Agnostic Federated MRI Reconstruction

### v1

@@ 3.2.2 MR Image Synthesis with the GAT Prior
To synthesize multi-site MR images via the GAT prior, the autoregressive transformer and VAE decoder modules are utilized. The transformer module generates a random set of multi-scale discrete token maps, sequentially across growing spatial scales based on the autoregressive mapping in Eq. 20. To introduce stochasticity during inference, however, nucleus sampling [62] is employed instead of the maximum-probability sampling procedure in Eqs. 18-19. Accordingly, the logit values across $V$ codebook vectors are ordered separately for each token in $\hat{\mathbf{f}}_{s}$, the top-5 percentile of values are subjected to softmax to obtain a probability mass function, and the predicted token is randomly drawn from this function:
@@ 3.2.3 Two-Tier Training Procedure for FedGAT
Training of the GAT prior: FedGAT decentrally trains a global, site-prompted GAT prior that captures the distribution of multi-site MR images, which is later used to generate synthetic MR images from desired sites. The proposed GAT prior embodies VAE and autoregressive transformer modules as described in Section 3.2.1. VAE aims to encode input MR images onto a compact latent space and then to decode them back faithfully. Accordingly, a composite loss function is used for the VAE modules [59]:
@@ 3 $\triangleright$ pre-train on local dataset
Training of reconstruction models: Following the first tier, each individual site trains a site-specific reconstruction model $H^{k}_{\phi^{k}}$ of preferred architecture, with the aid of the global GAT prior (Alg. 2). To attain a refined balance between within-site and cross-site reconstruction performances, the reconstruction model is built by pre-training on the local dataset followed by fine-tuning on a hybrid dataset containing both local and synthetic datasets. For pre-training, we employ a conventional mean-squared error loss [14]:
@@ 4.2 Competing Methods
FedGAT was comparatively demonstrated against non-FL benchmarks (Central, Single) along with distillation-based (FedMD, FedDF) and generative (FedGIMP, FedDDA) FL baselines.
@@
FedDF: A distillation-based FL baseline was considered that supports model-heterogeneous settings via an auxiliary public dataset [66]. To adopt FedDF for MRI reconstruction, distillation between site-specific reconstruction models and a global model were performed based on predictions of reconstructed images, instead of class logits as in the original implementation for classification tasks [66]. At individual sites, site-specific reconstruction models were initialized via distillation from the global model (taken as MoDL-5), and then trained along with the copy of the global model using a proximal loss term [67]. Training was performed on a 840-sample mixture of the local data with independent auxiliary data downloaded from a public repository [68]. On the server, copies of the global model locally trained at individual sites were aggregated via distillation on a 840-sample set of auxiliary data.
@@
FedDDA: A generative FL baseline that decentrally trains a global prior based on an unconditional diffusion model to synthesize images [50]. A site index was included in FedDDA for controllable generation of multi-site MR images [42]. FedDDA was imeplemented using the same two-tier training strategy as in FedGAT to support model-heterogeneous settings, where site-specific reconstruction models were trained using a hybrid dataset containing both local MRI data and synthetic data from remaining sites generated by FedDDA.
@@
Table 1: Reconstruction performance for single-coil datasets. Each site employs an entirely distinct model type, with Site $k$ (dataset-type) denoting the site index, the local dataset, and the type of reconstruction model. Metrics are listed as mean±$\pm$±std across the test set, for within-site (upper panel) and across-site (lower panel) reconstructions at R=4x-8x. Boldface indicates the top-performing FL method in each task. Benchmark methods for within-site reconstruction (Single), and for cross-site reconstruction (Central) are underlined.
Within-site Site 1 (BraTS-MoDL) Site 2 (fastMRI-rGAN) Site 3 (MIDAS-D5C5)
reconstruction PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
non-FL Single 4x 48.06$\pm$3.33 99.72$\pm$0.11 35.83$\pm$2.51 90.64$\pm$3.87 29.04$\pm$1.84 76.77$\pm$4.08
8x 42.46$\pm$3.15 99.03$\pm$0.45 33.18$\pm$2.26 86.58$\pm$4.64 27.79$\pm$2.05 74.20$\pm$3.97
Central 4x 46.12$\pm$2.55 99.46$\pm$0.20 36.03$\pm$2.51 91.23$\pm$3.63 28.89$\pm$1.89 75.11$\pm$4.07
8x 42.16$\pm$2.72 98.65$\pm$0.58 33.18$\pm$2.53 87.40$\pm$4.44 27.34$\pm$2.17 73.87$\pm$3.98
FL methods FedDF 4x 44.42$\pm$3.76 99.34$\pm$0.28 32.88$\pm$3.22 87.95$\pm$4.41 27.62$\pm$1.90 61.23$\pm$4.32
8x 40.06$\pm$3.23 98.23$\pm$0.77 30.78$\pm$2.73 83.45$\pm$5.51 27.48$\pm$1.93 59.62$\pm$4.93
FedMD 4x 44.68$\pm$3.72 99.40$\pm$0.24 34.24$\pm$3.06 88.48$\pm$4.13 28.95$\pm$1.92 62.22$\pm$5.25
8x 40.32$\pm$3.22 98.29$\pm$0.75 32.96$\pm$2.56 86.40$\pm$4.66 27.79$\pm$2.07 60.19$\pm$5.07
FedGIMP 4x 47.61$\pm$2.86 99.63$\pm$0.15 35.45$\pm$2.48 90.27$\pm$3.99 28.68$\pm$1.90 74.91$\pm$4.42
8x 42.14$\pm$2.71 98.62$\pm$0.54 32.74$\pm$2.40 85.86$\pm$4.75 26.82$\pm$2.17 73.96$\pm$4.08
FedDDA 4x 47.50$\pm$2.81 99.61$\pm$0.15 35.67$\pm$2.51 90.30$\pm$3.94 28.61$\pm$1.79 75.67$\pm$4.15
8x 41.76$\pm$2.90 98.47$\pm$0.63 32.79$\pm$2.45 86.02$\pm$4.84 27.10$\pm$2.18 73.08$\pm$4.07
FedGAT 4x 47.63$\pm$2.65 99.64$\pm$0.14 35.89$\pm$2.50 90.69$\pm$3.86 31.88$\pm$2.03 77.48$\pm$4.94
8x 41.60$\pm$2.85 98.78$\pm$0.54 33.23$\pm$2.36 86.45$\pm$0.67 29.62$\pm$2.11 74.88$\pm$4.29
Across-site Site 1 (BraTS-MoDL) Site 2 (fastMRI-rGAN) Site 3 (MIDAS-D5C5)
reconstruction PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
non-FL Single 4x 34.45$\pm$3.57 91.44$\pm$4.05 33.02$\pm$2.30 66.49$\pm$7.26 29.67$\pm$2.28 75.67$\pm$6.93
8x 30.48$\pm$3.20 85.48$\pm$5.37 30.08$\pm$2.20 59.14$\pm$6.78 28.38$\pm$2.26 73.64$\pm$6.55
Central 4x 36.48$\pm$2.89 94.78$\pm$1.91 33.53$\pm$2.31 82.93$\pm$4.07 30.68$\pm$2.47 77.41$\pm$6.99
8x 32.79$\pm$2.81 92.16$\pm$2.62 30.78$\pm$2.25 78.75$\pm$4.67 29.93$\pm$2.20 75.86$\pm$6.95
FL methods FedDF 4x 34.40$\pm$3.12 93.12$\pm$3.00 33.14$\pm$2.61 71.59$\pm$6.74 29.23$\pm$2.38 72.76$\pm$7.12
8x 30.91$\pm$2.87 90.33$\pm$3.36 30.89$\pm$2.22 64.17$\pm$7.14 28.37$\pm$2.63 72.45$\pm$8.43
FedMD 4x 34.61$\pm$3.13 93.42$\pm$2.75 33.11$\pm$2.26 68.04$\pm$6.84 30.03$\pm$2.41 72.31$\pm$7.93
8x 31.12$\pm$2.83 90.63$\pm$3.13 30.47$\pm$2.20 61.79$\pm$7.06 29.05$\pm$2.57 71.80$\pm$7.84
FedGIMP 4x 36.42$\pm$3.17 94.32$\pm$2.49 33.05$\pm$2.29 67.50$\pm$7.01 30.30$\pm$2.33 76.19$\pm$6.56
8x 31.88$\pm$2.87 90.16$\pm$3.78 29.78$\pm$2.16 57.80$\pm$6.39 28.31$\pm$2.16 71.75$\pm$6.10
FedDDA 4x 36.19$\pm$3.13 94.41$\pm$2.24 33.11$\pm$2.21 67.78$\pm$7.00 28.59$\pm$2.07 78.13$\pm$5.12
8x 31.79$\pm$2.86 90.75$\pm$3.07 30.15$\pm$2.16 59.18$\pm$6.93 27.73$\pm$2.13 77.92$\pm$5.11
FedGAT 4x 36.71$\pm$3.01 95.02$\pm$1.96 34.50$\pm$2.36 75.69$\pm$5.76 30.77$\pm$2.35 78.62$\pm$6.41
8x 32.02$\pm$2.95 91.53$\pm$3.07 31.47$\pm$2.49 71.84$\pm$6.47 29.72$\pm$2.38 78.08$\pm$6.86
@@
Table 2: Within-site and cross-site reconstruction performance for multi-coil datasets (f.knee denotes fastMRI-knee, f.brain denotes fastMRI-brain, UMRAM denotes the in-house dataset), where each site employs an entirely different model type.
Within-site Site 1 (f.knee-MoDL) Site 2 (f.brain-rGAN) Site 3 (UMRAM-D5C5)
reconstruction PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
non-FL Single 4x 33.78$\pm$2.27 93.37$\pm$1.52 36.97$\pm$3.07 97.32$\pm$1.17 31.81$\pm$1.73 90.10$\pm$2.57
8x 29.39$\pm$2.06 83.90$\pm$3.08 33.08$\pm$3.39 94.54$\pm$2.26 28.77$\pm$1.71 85.29$\pm$3.64
Central 4x 33.69$\pm$2.32 93.20$\pm$1.63 36.93$\pm$3.22 97.32$\pm$1.19 31.02$\pm$1.60 88.49$\pm$2.56
8x 29.39$\pm$2.19 83.62$\pm$3.09 33.79$\pm$3.05 94.59$\pm$2.13 27.39$\pm$1.68 82.72$\pm$3.04
FL methods FedDF 4x 32.79$\pm$2.19 90.40$\pm$2.45 33.95$\pm$2.88 96.30$\pm$1.39 30.34$\pm$1.55 88.99$\pm$2.53
8x 27.70$\pm$2.33 78.94$\pm$4.57 28.80$\pm$2.82 90.60$\pm$3.14 27.05$\pm$1.72 83.94$\pm$3.40
FedMD 4x 32.48$\pm$2.24 90.57$\pm$2.21 34.07$\pm$2.83 96.36$\pm$1.36 30.10$\pm$1.61 88.73$\pm$2.52
8x 26.99$\pm$2.47 78.49$\pm$4.51 29.08$\pm$3.15 90.16$\pm$3.23 27.13$\pm$1.64 84.18$\pm$3.31
FedGIMP 4x 33.78$\pm$2.27 93.41$\pm$1.48 31.72$\pm$3.85 94.25$\pm$4.32 30.11$\pm$2.08 88.56$\pm$2.43
8x 28.97$\pm$2.12 83.68$\pm$3.07 28.78$\pm$3.10 90.71$\pm$5.21 26.26$\pm$2.68 82.69$\pm$3.52
FedDDA 4x 32.53$\pm$2.33 91.75$\pm$1.82 32.69$\pm$3.46 94.71$\pm$3.44 24.47$\pm$2.57 81.20$\pm$3.47
8x 28.09$\pm$2.22 81.90$\pm$2.93 29.63$\pm$3.28 90.85$\pm$5.46 18.38$\pm$1.97 68.73$\pm$5.73
FedGAT 4x 34.10$\pm$2.24 97.32$\pm$1.67 35.86$\pm$2.78 96.85$\pm$1.19 31.94$\pm$1.70 90.21$\pm$2.51
8x 29.49$\pm$2.21 83.73$\pm$3.07 31.48$\pm$2.34 92.89$\pm$2.38 28.28$\pm$1.89 85.20$\pm$3.58
Across-site Site 1 (f.knee-MoDL) Site 2 (f.brain-rGAN) Site 3 (UMRAM-D5C5)
reconstruction PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
non-FL Single 4x 38.40$\pm$3.26 97.32$\pm$1.67 31.17$\pm$3.22 93.03$\pm$2.57 26.18$\pm$2.62 79.42$\pm$5.73
8x 32.39$\pm$3.04 92.63$\pm$2.94 26.82$\pm$3.15 85.13$\pm$5.06 22.64$\pm$2.30 70.31$\pm$6.85
Central 4x 39.70$\pm$3.36 97.62$\pm$1.70 37.41$\pm$2.13 96.27$\pm$1.29 27.52$\pm$2.49 81.63$\pm$5.35
8x 34.33$\pm$3.12 93.86$\pm$3.04 33.98$\pm$2.11 91.89$\pm$2.82 25.09$\pm$2.47 74.38$\pm$6.46
FL methods FedDF 4x 38.03$\pm$2.99 96.54$\pm$1.41 32.30$\pm$2.93 92.69$\pm$1.76 26.68$\pm$2.55 81.09$\pm$5.31
8x 32.05$\pm$2.80 92.25$\pm$3.12 27.68$\pm$2.80 85.12$\pm$4.50 24.04$\pm$2.48 70.60$\pm$6.94
FedMD 4x 37.77$\pm$3.07 96.53$\pm$1.59 32.62$\pm$2.94 92.69$\pm$1.69 26.39$\pm$2.50 80.99$\pm$5.26
8x 31.35$\pm$2.91 91.22$\pm$3.24 28.03$\pm$2.50 84.44$\pm$4.68 23.99$\pm$2.43 70.13$\pm$6.61
FedGIMP 4x 37.87$\pm$3.13 97.23$\pm$1.50 26.57$\pm$3.13 85.71$\pm$5.69 26.00$\pm$2.49 80.03$\pm$5.68
8x 32.01$\pm$2.98 92.52$\pm$2.91 22.85$\pm$2.93 75.96$\pm$8.28 22.20$\pm$2.29 67.89$\pm$8.82
FedDDA 4x 35.54$\pm$3.34 94.49$\pm$2.44 27.04$\pm$3.51 84.62$\pm$4.68 22.34$\pm$2.33 72.91$\pm$6.55
8x 30.15$\pm$3.57 88.74$\pm$3.98 23.98$\pm$3.40 75.66$\pm$5.86 19.21$\pm$2.32 60.73$\pm$9.18
FedGAT 4x 39.25$\pm$3.09 97.50$\pm$1.52 33.95$\pm$2.34 93.45$\pm$1.95 27.17$\pm$2.58 81.24$\pm$5.63
8x 33.52$\pm$2.98 92.90$\pm$3.11 29.20$\pm$2.28 86.58$\pm$4.23 24.45$\pm$2.55 77.63$\pm$8.77

### v2

@@ 3.2.2 MR Image Synthesis with the GAT Prior
To synthesize multi-site MR images via the GAT prior, the autoregressive transformer and VAE decoder modules are utilized. The transformer module generates a random set of multi-scale discrete token maps, sequentially across growing spatial scales based on the autoregressive mapping in Eq. 23. To introduce stochasticity during inference, however, nucleus sampling [71] is employed instead of the maximum-probability sampling procedure in Eqs. 21-22. Accordingly, the logit values across $V$ codebook vectors are ordered separately for each token in $\hat{\mathbf{f}}_{s}$, the top-5 percentile of values are subjected to softmax to obtain a probability mass function, and the predicted token is randomly drawn from this function:
@@ 3.2.3 Two-Tier Training Procedure for FedGAT
Tier 1 - Training of the GAT prior: FedGAT decentrally trains a global, site-prompted GAT prior that captures the distribution of multi-site MR images, which is later used to generate synthetic MR images from desired sites. The proposed GAT prior embodies VAE and autoregressive transformer modules as described in Section 3.2.1. VAE aims to encode input MR images onto a compact latent space and then to decode them back faithfully. Accordingly, a composite loss function is used for the VAE modules [68]:
@@ 3 $\triangleright$ pre-train on local dataset
Tier 2 - Training of reconstruction models: Following the first tier, each individual site trains a site-specific reconstruction model $H^{k}_{\phi^{k}}$ of preferred architecture, with the aid of the global GAT prior (Alg. 2). To attain a refined balance between within-site and cross-site reconstruction performances, the reconstruction model is built by pre-training on the local dataset followed by fine-tuning on a hybrid dataset containing both local and synthetic datasets. For pre-training, we employ a conventional mean-squared error loss [13]:
@@ 4.2 Competing Methods
FedGAT was comparatively demonstrated against non-FL benchmarks (Central, Single) along with distillation-based (FedMD, FedDF) and generative (FedGIMP, FedDDA) FL baselines. Since previous FL methods for MRI reconstruction were not specifically designed to handle model-heterogeneous settings, we either adopted model-agnostic baselines from non-reconstruction tasks (FedMF, FedDG, FedDDA), or modified reconstruction baselines to accommodate model heterogeneity when feasible (FedGIMP).
@@
FedDF: A distillation-based FL baseline was considered that supports model-heterogeneous settings via an auxiliary public dataset [75]. To adopt FedDF for MRI reconstruction, distillations between site-specific reconstruction models and a global model were performed based on predictions of reconstructed images, instead of class logits as in the original implementation for classification tasks [75]. At individual sites, site-specific reconstruction models were initialized via distillation from the global model (taken as MoDL5), and then trained along with the copy of the global model using a proximal loss term [76]. Training was performed on a 840-sample mixture of the local data with independent auxiliary data downloaded from a public repository [77]. On the server, copies of the global model locally trained at individual sites were aggregated via distillation on a 840-sample set of the auxiliary data.
@@
FedDDA: A generative FL baseline that decentrally trains a global prior based on an unconditional diffusion model to synthesize images [59]. A site index was included in FedDDA for controllable generation of multi-site MR images [51]. FedDDA was implemented using the same two-tier training strategy as in FedGAT to support model-heterogeneous settings, where site-specific reconstruction models were trained using a hybrid dataset containing both local MRI data and synthetic data from remaining sites generated by FedDDA.
@@
Table 1: Reconstruction performance for coil-combined datasets (Site 1: BraTS, Site 2: fastMRI, Site 3: MIDAS). Site $k$ (type) denotes site index and type of reconstruction model. Metrics are listed as mean±$\pm$±std across the test set, for within-site (upper panel) and across-site (lower panel) reconstructions at R=4x-8x. Boldface indicates the top-performing FL method in each task. Benchmark methods for within-site reconstruction (Single), and for cross-site reconstruction (Central) are underlined.
Within-site Site 1 (MoDL) Site 2 (rGAN) Site 3 (D5C5) Average
[…]
non-FL Single 4x 48.1$\pm$3.3 99.7$\pm$0.1 35.8$\pm$2.5 90.6$\pm$3.9 29.0$\pm$1.8 76.8$\pm$4.1 37.6 89.0
8x 42.5$\pm$3.2 99.0$\pm$0.5 33.2$\pm$2.3 86.6$\pm$4.6 27.8$\pm$2.1 74.2$\pm$4.0 34.5 86.6
Central 4x 46.1$\pm$2.6 99.5$\pm$0.2 36.0$\pm$2.5 91.2$\pm$3.6 28.9$\pm$1.9 75.1$\pm$4.1 37.0 88.6
8x 42.2$\pm$2.7 98.7$\pm$0.6 33.2$\pm$2.5 87.4$\pm$4.4 27.3$\pm$2.2 73.9$\pm$4.0 34.2 86.7
FL methods FedDF 4x 44.4$\pm$3.8 99.3$\pm$0.3 32.9$\pm$3.2 88.0$\pm$4.4 27.6$\pm$1.9 61.2$\pm$4.3 35.0 82.8
8x 40.1$\pm$3.2 98.2$\pm$0.8 30.8$\pm$2.7 83.5$\pm$5.5 27.5$\pm$1.9 59.6$\pm$4.9 32.8 80.4
FedMD 4x 44.7$\pm$3.7 99.4$\pm$0.2 34.2$\pm$3.1 88.5$\pm$4.1 29.0$\pm$1.9 62.2$\pm$5.3 36.0 83.4
8x 40.3$\pm$3.2 98.3$\pm$0.8 33.0$\pm$2.6 86.4$\pm$4.7 27.8$\pm$2.1 60.2$\pm$5.1 33.7 81.6
FedGIMP 4x 47.6$\pm$2.9 99.6$\pm$0.2 35.5$\pm$2.5 90.3$\pm$4.0 28.7$\pm$1.9 74.9$\pm$4.4 37.3 88.3
8x 42.1$\pm$2.7 98.6$\pm$0.5 32.7$\pm$2.4 85.9$\pm$4.8 26.8$\pm$2.2 74.0$\pm$4.1 33.9 86.2
FedDDA 4x 47.5$\pm$2.8 99.6$\pm$0.2 35.7$\pm$2.5 90.3$\pm$3.9 28.6$\pm$1.8 75.7$\pm$4.2 37.3 88.5
8x 41.8$\pm$2.9 98.5$\pm$0.6 32.8$\pm$2.5 86.0$\pm$4.8 27.1$\pm$2.2 73.1$\pm$4.1 33.9 85.9
FedGAT 4x 47.9$\pm$2.7 99.7$\pm$0.1 35.9$\pm$2.5 90.7$\pm$3.9 31.9$\pm$2.0 77.5$\pm$4.9 38.6 89.3
8x 41.6$\pm$2.9 98.7$\pm$0.6 33.2$\pm$2.4 86.5$\pm$0.7 29.6$\pm$2.1 74.9$\pm$4.3 34.8 86.7
Across-site Site 1 (MoDL) Site 2 (rGAN) Site 3 (D5C5) Average
[…]
non-FL Single 4x 34.5$\pm$3.6 91.4$\pm$4.1 33.0$\pm$2.3 66.5$\pm$7.3 29.7$\pm$2.3 75.7$\pm$6.9 32.4 77.9
8x 30.5$\pm$3.2 85.5$\pm$5.4 30.1$\pm$2.2 59.1$\pm$6.8 28.4$\pm$2.3 73.6$\pm$6.6 29.7 72.7
Central 4x 36.5$\pm$2.9 94.8$\pm$1.9 33.5$\pm$2.3 82.9$\pm$4.1 30.7$\pm$2.5 77.4$\pm$7.0 33.6 85.0
8x 32.8$\pm$2.8 92.2$\pm$2.6 30.8$\pm$2.3 78.8$\pm$4.7 29.9$\pm$2.2 75.9$\pm$7.0 31.2 82.3
FL methods FedDF 4x 34.4$\pm$3.1 93.1$\pm$3.0 33.1$\pm$2.6 71.6$\pm$6.7 29.2$\pm$2.4 72.8$\pm$7.1 32.2 79.2
8x 30.9$\pm$2.9 90.3$\pm$3.4 30.9$\pm$2.2 64.2$\pm$7.1 28.4$\pm$2.6 72.5$\pm$8.4 30.1 75.7
FedMD 4x 34.6$\pm$3.1 93.4$\pm$2.8 33.1$\pm$2.3 68.0$\pm$6.8 30.0$\pm$2.4 72.3$\pm$7.9 32.6 77.9
8x 31.1$\pm$2.8 90.6$\pm$3.1 30.5$\pm$2.2 61.8$\pm$7.1 29.1$\pm$2.6 71.8$\pm$7.8 30.2 74.7
FedGIMP 4x 36.4$\pm$3.2 94.3$\pm$2.5 33.1$\pm$2.3 67.5$\pm$7.0 30.3$\pm$2.3 76.2$\pm$6.6 33.3 79.3
8x 31.9$\pm$2.9 90.2$\pm$3.8 29.8$\pm$2.2 57.8$\pm$6.4 28.3$\pm$2.2 71.8$\pm$6.1 30.0 73.3
FedDDA 4x 36.2$\pm$3.1 94.4$\pm$2.2 33.1$\pm$2.2 67.8$\pm$7.0 28.6$\pm$2.1 78.1$\pm$5.1 32.6 80.1
8x 31.8$\pm$2.9 90.8$\pm$3.1 30.2$\pm$2.2 59.2$\pm$6.9 27.7$\pm$2.1 77.9$\pm$5.1 29.9 76.0
FedGAT 4x 36.7$\pm$4.3 95.0$\pm$2.0 34.5$\pm$2.4 75.7$\pm$5.8 30.8$\pm$2.4 78.6$\pm$6.4 34.0 83.1
8x 32.1$\pm$4.0 91.5$\pm$3.1 31.5$\pm$2.5 71.8$\pm$6.5 29.7$\pm$2.4 78.1$\pm$6.9 31.1 80.5
@@
Table 2: Within-site and cross-site reconstruction performance for multi-coil datasets (Site 1: fastMRI-knee, Site 2: fastMRI-brain, Site 3: UMRAM). Each site employs a distinct model type.
Within-site Site 1 (MoDL) Site 2 (rGAN) Site 3 (D5C5) Average
[…]
non-FL Single 4x 33.8$\pm$2.3 93.4$\pm$1.5 37.0$\pm$3.1 97.3$\pm$1.2 31.8$\pm$1.7 90.1$\pm$2.6 34.2 93.6
8x 29.4$\pm$2.1 83.9$\pm$3.1 33.1$\pm$3.4 94.5$\pm$2.3 28.8$\pm$1.7 85.3$\pm$3.6 30.4 87.9
Central 4x 33.7$\pm$2.3 93.2$\pm$1.6 36.9$\pm$3.2 97.3$\pm$1.2 31.0$\pm$1.6 88.5$\pm$2.6 33.9 93.0
8x 29.4$\pm$2.2 83.6$\pm$3.1 33.8$\pm$3.1 94.6$\pm$2.1 27.4$\pm$1.7 82.7$\pm$3.0 30.2 87.0
FL methods FedDF 4x 32.8$\pm$2.2 90.4$\pm$2.5 34.0$\pm$2.9 96.3$\pm$1.4 30.3$\pm$1.6 89.0$\pm$2.5 32.4 91.9
8x 27.7$\pm$2.3 78.9$\pm$4.6 28.8$\pm$2.8 90.6$\pm$3.1 27.1$\pm$1.7 83.9$\pm$3.4 27.9 84.5
FedMD 4x 32.5$\pm$2.2 90.6$\pm$2.2 34.1$\pm$2.8 96.4$\pm$1.4 30.1$\pm$1.6 88.7$\pm$2.5 32.2 91.9
8x 27.0$\pm$2.5 78.5$\pm$4.5 29.1$\pm$3.2 90.2$\pm$3.2 27.1$\pm$1.6 84.2$\pm$3.3 27.7 84.3
FedGIMP 4x 33.2$\pm$2.3 93.1$\pm$1.5 31.7$\pm$3.9 94.3$\pm$4.3 30.1$\pm$2.1 88.6$\pm$2.4 31.7 92.0
8x 29.2$\pm$2.1 83.6$\pm$3.1 28.8$\pm$3.1 90.7$\pm$5.2 26.3$\pm$2.7 82.7$\pm$3.5 28.1 85.7
FedDDA 4x 32.5$\pm$2.3 91.8$\pm$1.8 32.7$\pm$3.5 94.7$\pm$3.4 24.5$\pm$2.6 81.2$\pm$3.5 29.9 89.2
8x 28.1$\pm$2.2 81.9$\pm$2.9 29.6$\pm$3.3 90.9$\pm$5.5 18.4$\pm$2.0 68.7$\pm$5.7 25.4 80.5
FedGAT 4x 34.1$\pm$2.2 97.3$\pm$1.7 35.9$\pm$2.8 96.9$\pm$1.2 31.9$\pm$1.7 90.2$\pm$2.5 34.0 94.8
8x 29.5$\pm$2.2 83.7$\pm$3.1 31.5$\pm$2.3 92.9$\pm$2.4 28.3$\pm$1.9 85.2$\pm$3.6 29.8 87.3
Across-site Site 1 (MoDL) Site 2 (rGAN) Site 3 (D5C5) Average
[…]
non-FL Single 4x 38.4$\pm$3.3 97.3$\pm$1.7 31.2$\pm$3.2 93.0$\pm$2.6 26.2$\pm$2.6 79.4$\pm$5.7 31.9 89.9
8x 32.4$\pm$3.0 92.6$\pm$2.9 26.8$\pm$3.2 85.1$\pm$5.1 22.6$\pm$2.3 70.3$\pm$6.9 27.3 82.7
Central 4x 39.7$\pm$3.4 97.6$\pm$1.7 37.4$\pm$2.1 96.3$\pm$1.3 27.5$\pm$2.5 81.6$\pm$5.4 34.9 91.8
8x 34.3$\pm$3.1 93.9$\pm$3.0 34.0$\pm$2.1 91.9$\pm$2.8 25.1$\pm$2.5 74.4$\pm$6.5 31.1 86.7
FL methods FedDF 4x 38.0$\pm$3.0 96.5$\pm$1.4 32.3$\pm$2.9 92.7$\pm$1.8 26.7$\pm$2.6 81.1$\pm$5.3 32.3 90.1
8x 32.1$\pm$2.8 92.3$\pm$3.1 27.7$\pm$2.8 85.1$\pm$4.5 24.0$\pm$2.5 70.6$\pm$6.9 27.9 82.6
FedMD 4x 37.8$\pm$3.1 96.5$\pm$1.6 32.6$\pm$2.9 92.7$\pm$1.7 26.4$\pm$2.5 81.0$\pm$5.3 32.3 90.1
8x 31.4$\pm$2.9 91.2$\pm$3.2 28.0$\pm$2.5 84.4$\pm$4.7 24.0$\pm$2.4 70.1$\pm$6.6 27.8 81.9
FedGIMP 4x 37.6$\pm$3.2 97.2$\pm$1.6 26.6$\pm$3.1 85.7$\pm$5.7 26.0$\pm$2.5 80.0$\pm$5.7 30.1 87.6
8x 32.1$\pm$3.0 92.5$\pm$2.9 22.9$\pm$2.9 76.0$\pm$8.3 22.2$\pm$2.3 67.9$\pm$8.8 25.7 78.8
FedDDA 4x 35.5$\pm$3.3 94.5$\pm$2.4 27.0$\pm$3.5 84.6$\pm$4.7 22.3$\pm$2.3 72.9$\pm$6.6 28.3 84.0
8x 30.2$\pm$3.6 88.7$\pm$4.0 24.0$\pm$3.4 75.7$\pm$5.9 19.2$\pm$2.3 60.7$\pm$9.2 24.5 75.0
FedGAT 4x 39.3$\pm$3.1 97.5$\pm$1.5 34.0$\pm$2.3 93.5$\pm$2.0 27.2$\pm$2.6 81.2$\pm$5.6 33.5 90.7
8x 33.5$\pm$3.0 92.9$\pm$3.1 29.2$\pm$2.3 86.6$\pm$4.2 24.5$\pm$2.6 77.6$\pm$8.8 29.1 85.7